OpenStudy (anonymous):

FAN AND MEDAL AWARDED Evaluate the expression, using the given value of the variable. z ÷ 3 – 5z + 4(z – 2) when z = 12 A. –98 B. –96 C. –16 D. –14

OpenStudy (aravindg):

Ok use BODMAS rule and substitute x=12 in the equation

OpenStudy (aravindg):

@alahrichi1

OpenStudy (anonymous):

I did that and got -5

OpenStudy (aravindg):

try again?

OpenStudy (aravindg):

can u please show ur step then i can know where you are making mistake

OpenStudy (aravindg):

@alahrichi1 ?

OpenStudy (anonymous):

=12+3-5*12+4(12-2) =12+3-5*12+4*10 =12+3-60+4*10 =12+3-60+40 =15-100 =-85

OpenStudy (aravindg):

ok ur first step is right

OpenStudy (aravindg):

you considered the brackets and got 10

OpenStudy (anonymous):

yes..

OpenStudy (aravindg):

and note it is z/3 not z+3 !!

OpenStudy (anonymous):

OMG!!!!!!!!!!!! thanks sooo much!! thats where my mistake was!!!

OpenStudy (aravindg):

=12/3-5*12+4*10 =12/3-5*12+40 now do division

OpenStudy (aravindg):

yw :)

OpenStudy (anonymous):

ok 1 sec

OpenStudy (anonymous):

-96

OpenStudy (aravindg):

try again?

OpenStudy (aravindg):

show me ur step after =12/3-5*12+4*10 =12/3-5*12+40

OpenStudy (anonymous):

=4-60+40 =-96 :D

OpenStudy (aravindg):

-60+40=??

OpenStudy (aravindg):

reply pls i gotta sleep after finishing this !!

OpenStudy (aravindg):

@alahrichi1 .......

OpenStudy (anonymous):

-20

OpenStudy (aravindg):

-20+4=?

OpenStudy (anonymous):

no bro, not -60+40 its 4 minus 60

OpenStudy (aravindg):

both are same :)

OpenStudy (aravindg):

subtraction is actually a form of addition only :)

OpenStudy (aravindg):

whatevr way you do the answer will be same

OpenStudy (aravindg):

4-60=?

OpenStudy (aravindg):

please reply fast ?

OpenStudy (anonymous):

-56

OpenStudy (aravindg):

-56+40 =?

OpenStudy (anonymous):

-16

OpenStudy (aravindg):

:D you done it !!!

OpenStudy (aravindg):

if you did other way also you will get -16 see =4-60+40 =4-20 =-16

OpenStudy (aravindg):

beauty of mathematics !!

OpenStudy (anonymous):

omg dude u are a legend

OpenStudy (aravindg):

:D thanks ..nw i can sleep :)

OpenStudy (anonymous):

good nite

OpenStudy (aravindg):

thanks hav good day/night :)

OpenStudy (anonymous):

u too